# Nano Drone-based Indoor Crime Scene Analysis*

### v1

@@ III METHOD
Then, we constructed proof-of-concepts. For hardware, a DJI Ryze Tello drone was augmented to be able to see both below and in front, by popping out its side camera to point downwards, and adding a 3D printed stand on top of its chassis to carry an ESP32-Cam, powered by a lithium polymer battery. The Ryze Tello drone is inexpensive, easily programmed, and has a large payload (> 60 g); ESP32-Cam is also inexpensive and easily programmed to wirelessly stream SVGA@30fps from its 2 Megapixel OV2640 camera to a python client. We also set up some software on an external laptop, including OpenCV181818opencv.org and YOLO 11.191919docs.ultralytics.com/models/yolo11/ Fig. 2 and Fig. 3 visually depict some of the challenges and successes we encountered using this setup, which are detailed below.
Figure 2: Examples of initial challenges: (a) getting caught, (b) markers not being detected, and (c) false detections in poor illumination
@@
III-A Capability 1: Situation. Access via a partially open window
@@
Rapid access is a crucial reason for using drones, and humans cannot be expected to be on scene to let them in, so drones could seek to gain entry in various ways, including via windows. In some cases, however, a window might be partially open, such that acrobatic entry could be difficult, and lock-picking unnecessary, but how to handle such cases was unclear. Although hardware is continually advancing, clearly it could be difficult currently for a nano drone to force open a tough window. For example, Yeong et al. estimate that opening a door requires 35 N [37]–whereas, based on Newton’s second law of motion (F=ma), a drone like a Ryze Tello with approximate mass 100 g and acceleration 5m/s2 would have only 0.5 N of thrust force. While there are larger drones like DJI Matrice and first person view (FPV) racing drones with high acceleration, safety was also a concern.
@@
Thus, our goal was to explore how to safely get started with testing ideas. We chose to prepare an initial mock-up window using a cardboard box with plastic wrap in place of glass. Of the four basic kinds of window–rotating horizontally (casement) or vertically (awning), and translating horizontally (sliding) or vertically (double hung)–a common "casement" form was selected. Various opening mechanisms for the drone were also considered, from wedges to actuated tongs. Eventually, a simplified approach was selected by adding a plastic spike to the underside of the drone with putty.
@@
Initial trials with manually piloting the drone to open the window with the spike indicated substantial challenges with low force, downwash (inaccurate control), and propellers getting caught, with a performance of roughly only 20%. Therefore, we redesigned the test window to be lighter (91 g), without deep walls, and with paper in place of plastic wrap, and conducted 20 trials. The result was 75% successful performance. A main factor in the five failed attempts was the imperfect control due to the inexpensive drone used: for example, the drone sometimes came into contact with the window near the hinges where torque would be insufficient or was knocked outside of the area of the window.
III-B Capability 2: Task. Mapping evidence
@@
After gaining access, a drone should map the crime scene, calculating distances between samples and gathering evidence, which was unclear how to realize. Our approach consisted in setting the drone to fly twice, once to create a 3D map, and the second time to find evidence. To simplify our initial exploration, a small 2m x 2m mock-up murder crime scene was created using some photos of guns, bullet casings, and blood, and four ArUco markers to mark walls. The gun was placed on a table, and the other photos on the ground.
@@
First, a 3D map was initialized. The Ryze Tello drone took off and rotated to detect markers via its side camera (ESP32-Cam), whose video feed was streamed through WiFi to an external laptop. The first marker detected was set to be located at the origin relative to the x and z axes and at a height along the y axis inferred based on the drone’s height given by its bottom infrared range sensor. As the drone rotated, newly detected markers were sequentially localized with respect to the first marker. Once mapping was complete, the drone landed.
@@
Second, the drone took off again, looking for evidence within the 3D map, by moving between markers. The drone’s video feed from its downward facing camera was again relayed to the external laptop, which used YOLO 11 with a GPU to detect three classes of object (blood, guns, and bullet casings). Visual servoing was conducted to center the drone over the detected evidence. Then, information about evidence was added to the 3D map, based on the drone’s estimated location and altitude. Furthermore, the drone was set to land on the center of the detected evidence to mimic gathering a sample (e.g., harvesting blood via a swab on its underside). Finally, when the drone’s search ended and its map was completed, the distances between all detected objects was calculated and output.
@@
Initial trials indicated problems with ArUco markers being too small to be detected when the drone was far or too big when the drone was close. Some trials were conducted with multiple markers. Eventually, a 4x4 ChArUco Board was adopted, which combines a chess board pattern and multiple ArUco markers to enhance accuracy and handle partial occlusions.
@@
To evaluate the system, 20 trials were conducted. Accuracy of detecting objects was 85.0%. Reasons for some inaccuracy included illumination and lack of contrast between the bloodstain and the floor, as well as limited amounts of training images. The average discrepancy between actual distances and estimated distances was also found to be 2.395 cm (SD 0.496 cm). This error might have been due to challenges with the stability of the drone when capturing images, illumination, and the limited resolution of the drone’s cameras (ESP32-Cam), as a lower resolution was used to avoid lag. Additionally, we noted that YOLO worked at 24 fps, and trials took approximately three minutes.
@@
Once evidence has been documented, it should be analyzed. For example, BPA on transfer stains such as smears and swipes could indicate where a criminal went or if a body was moved, given that it could be hidden to delay or confuse investigators.
@@
For initial exploration, a simplified dataset was created using red carmine dye (cochineal- E120) on white paper. Body parts, continuity, and direction were varied. Continuity involved if the body part was constantly in contact with the paper ("legato", like a corpse being dragged) or set down at regular intervals (e.g., "staccato", like a shoe). Specifically, we walked over paper with dyed shoes and crawled on our hands to mimic a criminal or victim fleeing; and, dragged a hand, and touched down or dragged an amorphous blob using a soaked tissue, to mimic a corpse being moved. In total, twenty samples were created: five samples of each were obtained, which were viewed four times each by a rotating drone to point up, down, left, and right.
@@
A python program was written to calculate the perceived direction of motion using a simplified approach involving color picking and image moments, given that the centroid should be closer to the start position where there is more blood than the end position where little blood remains. First, red color was detected via hue (two ranges), then contours. Then a line was fitted to the centroids of contours. Finally, the position of the centroid of centroids along the line was used to estimate the direction the bloody object was moved with dot product and an adjustment for the quadrant. A codebook vector/class label was selected as the closest to the resulting angle, and the error between prediction and ground truth calculated.
@@
Initial trials showed various challenges, such as many small false contours using standard hue values for red, lines not being fit for long "legato" smears with a single contour, as well as some errors estimating direction with multiple contours, given that area was not taken into account. Therefore, the range for orange-red hue was removed, contour area was checked, code was written to handle base cases (e.g., if only one contour, which can happen if blood picking is good and an object is being dragged, moments are used to find a line based on the contour’s shape), and the midpoint of the detected blood pixels was used in place of the center of centroids to estimate direction.
@@
As a result, accuracy was 80%, with an average error of 50.3 degrees. Hand staccato, shoe staccato, hand legato were completely detected. However, the amorphous "tissue" stains seemed to be predicted at the level of random chance (50%). We noted that we could also not tell from observation ourselves which side was the start or end point. As well, the reason the average error was large was because of the four error cases, mostly again the "tissue" samples: three were 180 degrees wrong (in other words, the "line of motion" is estimated correctly in 19 of 20 cases, so that the error would be much less if the start and end are not required). Thus, the results indicated that some cases can be easier to infer than others.
@@ IV DISCUSSION
We presented an overview of indoor crime scene analysis (CSA) by a nano drone. First, we identified theoretical gaps, based on a "big picture" (rapid scoping) review. Second, practical challenges were explored by iteratively designing three prototype capabilities with performances of 75%, 85%, and 80%. While nano drones do not seem capable yet of, e.g., opening windows, skillfully gathering evidence, and analyzing complex blood patterns in the real world, we believe that our results suggest the feasibility of initial ideation and testing within a laboratory setting. A video summarizes this work.  [^20] https://youtu.be/An8TIGr_NRA
@@ IV-A Limitations and Future Work
Our approach is exploratory, involving rapid, low-fidelity prototyping; i.e., using red paint in place of blood and photos of guns instead of real firearms. Furthermore, wireless transmission of potentially sensitive data from drones to remote laptops for processing could present an security threat, allowing criminals to intercept and possibly alter communications between drone and humans. As well, the scope of this paper did not allow us to explore all of the identified gaps.
Future work should move farther from the laboratory into the real world. Stronger drones will autonomously fly in more accurate settings, onboard sensors like lidars will be used rather than markers for mapping, and real blood stain patterns will be analyzed. Processing should be preferably done by computers on the drone. Additionally, we have started to explore a broader range of capabilities we identified as potential gaps, conducting some development on each of our ideas. These additional scenarios are visualized in Fig. 4, and in the accompanying video.
Figure 4: (a) Revealing hidden blood stains, (b) documenting evidence without disturbance via a blimp, (c) using a flashlight to operate in the dark, and (d) detecting hidden people with thermal or radar.
In conclusion, while much work remains to be done to make this vision of nano drones helping to solve crimes a reality, we believe that the day is not too far off in which such technologies will be help preserve justice and impact society in a positive way.

### v2

@@ III METHOD
Then, we constructed proof-of-concepts. For hardware, a DJI Ryze Tello drone [27] was augmented to be able to see both below and in front, by popping out its side camera to point downwards, and adding a 3D printed stand on top of its chassis to carry an ESP32-Cam, powered by a lithium polymer battery, as shown in Fig. 2, The Ryze Tello drone is inexpensive, easily programmed, and has a large payload (> 60 g); ESP32-Cam is also inexpensive and easily programmed to wirelessly stream SVGA@30fps from its 2 Megapixel OV2640 camera [65]. We also set up some software on an external laptop, including OpenCV [66] and YOLO 11 [67]. An overview of our set-up for exploring the three capabilities is shown in Fig. 3. Also, Fig. 4 and Fig. 5 visually depict some of the challenges and successes encountered using this setup, which are detailed below.
Figure 2: Drone used for exploration, a modified Ryze Tello
Figure 3: Set-up for exploring each capability: (a) the drone rises and tries to move straight through a partially open "window", (b) the drone rises and moves between markers, pausing to "gather" detected evidence, and (c) the drone rises and rotates above a "blood smear", estimating direction.
Figure 4: Examples of initial challenges: (a) propellers getting caught, (b) single markers not being detected, and (c) false detections in low light
@@
III-A Capability 1: Situation. Pushing Open a Window
@@
Rapid access is a crucial reason for using drones, and humans cannot be expected to be on scene to let them in, so drones could seek to gain entry in various ways, including via windows. In some cases, a window might be partially open, such that acrobatic entry could be difficult, and lock-picking unnecessary, but how to handle such cases was unclear.
Yeong et al. (on page 4082, section IIIA, paragraph 3, line 8) estimated that opening a door requires 35 N [68]. Assuming similar forces required for windows and doors, a drone like the Drone Racing League (DRL) Racer3 with 71 N of thrust–possessing a hard shell to protect from crashes, weighing over 1kg and capable of accelerating to 36m/s in under a second–could potentially push open some partially opened windows [69]. However, safety is also a concern, since an error at high acceleration could lead to injuries. This suggested the usefulness of starting out with a nano drone for safely testing ideas.
@@
Although hardware is continually advancing, it could be difficult currently for a nano drone to force open a tough window. Based on Newton’s second law of motion (F=ma), a small drone with approximate mass 100 g and acceleration 5m/s2 would have only 0.5 N of thrust force. Thus, we chose to prepare an initial mock-up window using a cardboard box with plastic wrap in place of glass. Of the four basic kinds of window–rotating horizontally (casement) or vertically (awning), and translating horizontally (sliding) or vertically (double hung)–a common casement form was selected. Various opening mechanisms for the drone were also considered, from wedges to actuated tongs. Eventually, for initial testing, a plastic spike was attached to the underside of the drone with putty.
@@
Initial trials with manually piloting the drone to open the window with the spike indicated substantial challenges with low force, downwash (inaccurate control), and propellers getting caught despite protective guards, with a performance of roughly only 20%. Therefore, we redesigned the test window to be lighter (91 g), without deep walls (string was used to hang it from the ceiling and pull the bottom down via a weight on the floor), and with paper in place of plastic wrap. The revised set-up was checked by conducting 20 trials. The result was 75% successful performance, with the drone pushing through the partially opened window 15 times. A main factor in the five failed attempts was the imperfect control due to the inexpensive drone used: for example, the drone sometimes came into contact with the window near its hinges where torque would be insufficient or was knocked outside of the area of the window.
III-B Capability 2: Tasks. Mapping Evidence
@@
After gaining access, a drone should map the crime scene, calculating distances between samples and gathering evidence, which was unclear how to realize. To simplify our initial exploration, we set the drone to fly twice, once to create a 3D map, and the second time to find evidence. Also, a small 2 m x 2 m mock-up murder crime scene was created using some printed photos of a gun, blood, and bullet casings on A3 paper (29.7 cm width x 42.0 cm height), and four ArUco markers to mark walls. The photo of a gun was placed on a table, and the other photos on the ground.
@@
First, a 3D map was initialized. The Ryze Tello drone took off and rotated to detect markers via its side camera (ESP32-Cam), whose video feed was streamed through WiFi to an external laptop. The first marker detected was assigned to a location at the origin relative to the x and z axes and at a height along the y axis inferred based on the drone’s height, given by its bottom infrared range sensor. As the drone rotated, newly detected markers were sequentially localized with respect to the first marker. Once mapping was complete, the drone landed.
@@
Second, the drone took off again, looking for evidence within the 3D map, by moving between markers. The drone’s video feed from its downward facing camera was again relayed to the external laptop, which used YOLO 11 with a GPU to detect three classes of object (guns, blood, and bullet casings). Visual servoing was conducted to center the drone over the detected evidence. Then, information about evidence was added to the 3D map, based on the drone’s estimated location and altitude. Furthermore, the drone was set to land on the center of the detected evidence to mimic gathering a sample (e.g., harvesting blood via a swab on its underside). Finally, when the drone’s search ended and its map was completed, the distances between all detected objects was calculated and output.
@@
Initial trials indicated problems with ArUco markers being too small to be detected when the drone was far or too big when the drone was close. Eventually, a 4x4 ChArUco Board was adopted, which combines a chess board pattern and multiple ArUco markers to enhance accuracy and handle partial occlusions.
@@
To evaluate the system, 20 trials were conducted, which resulted in an object detection accuracy of 85.0%. Reasons for some inaccuracy included illumination and lack of contrast between the bloodstain photo and the floor, as well as limited amounts of training images. The average discrepancy between actual distances and estimated distances was also found to be 2.4 cm (SD: 0.5 cm). This error might have been due to challenges with the stability of the drone when capturing images, illumination, and the limited resolution of the drone’s cameras (ESP32-Cam), as a lower resolution was used to avoid lag. Additionally, we noted that YOLO worked at 24 fps, and trials took approximately three minutes.
As noted by Bucknell and Bassindale, however, a drone moving overhead can disturb trace evidence like fibers, hairs, or small particles [50]. We expect this effect to be strong here, given that the drone flies twice, and even lands on evidence to simulate gathering. Thus, to avoid damage, such an approach should only be used when the absence of trace evidence has been verified, or an investigator has decided that the risk is acceptable.
@@
Once evidence has been documented, it should be analyzed. For example, BPA on transfer stains such as smears and swipes could indicate where a criminal went or a body was moved (e.g., to delay or confuse investigators).
@@
For initial exploration, since we could not find a dataset that would fit the designated context, a simplified dataset was created using red paint on white paper as a substitute for human blood: viz., carmine dye (E120, from cochineal insects), on four white A4 sheets per sample (21.0 cm width x 118.8 cm length). Body parts, continuity, and direction were varied. Body parts included a hand, shoe, and an amorphous blob. Continuity involved if the body part was constantly in contact with the paper ("legato", like a corpse being dragged) or set down at regular intervals ("staccato", like a shoe). Specifically, we walked over the paper with a soaked shoe (European size 43) and crawled with one soaked hand (length: 19.5 cm) to mimic a criminal or victim fleeing; and, we dragged a soaked hand, and touched down or dragged an amorphous blob using a soaked tissue, to mimic a corpse being moved. In total, twenty samples were created: Our drone was programmed to automatically take off and rotate over each of the five samples, capturing video (320 x 240). Then, for each video, images were manually captured from four views: up, down, left, and right.
@@
Next, a Python program was written to calculate the perceived direction of motion, using a simplified approach involving color picking and image moments. The intuition was that the majority of the detected red color should be closer to the start position where there is more blood than the end position where little blood remains. First, red color was detected via hue (two ranges), then contours. Then a line was fitted to the centroids of contours. Finally, the position of the centroid of centroids along the line was used to estimate the direction the bloody object was moved with dot product and an adjustment for the quadrant. A closest codebook vector/class label was selected, and the error between prediction and ground truth calculated.
@@
Initial trials showed various challenges, such as many small false contours using standard hue values for red (possibly due to illumination and the dye used), lines not being fit for long "legato" smears with a single contour, as well as some errors estimating direction with multiple contours from assigning equal weight to large and tiny contours. Therefore, the range for red was adjusted (e.g., removing orange-red), contour area was checked, code was written to handle base cases (e.g., in cases with only one contour, which can happen if an object is being dragged, moments are used to find a line based on the contour’s shape), and the midpoint of the detected blood pixels was used in place of the center of centroids to estimate direction.
@@
As a result, accuracy was 80%, with an average error of 50.3°. Accuracy was imperfect due to a similar amount of red color being detected at the start and end for two samples: While hand staccato, shoe staccato, hand legato were completely detected, the amorphous "tissue" stains seemed to be predicted at the level of random chance (50%), and we could also not tell from observation which side was the start or end point.
Likewise, the average error was large for the same reason, the amorphous tissue stains. Of the four error cases, three were off by 180°, since start and end points were reversed. We feel that this error is not a crucial problem for the following reasons: The "line of motion" is estimated correctly in 19 of 20 cases (95% accuracy). We don’t expect all transfer stains at crime scenes to be amorphous tissue stains; the error for correct cases was 26.2°, which seems reasonable. Even considering the challenging tissue stains, 50.3°is less than the error expected from a random guess, which would be 180°. As well, our goal is only to explore one possible simplified approach to give an initial estimation; the approach can be improved, and the initial estimate later verified by humans or robots. Thus, the results indicated that some cases can be easier to infer than others.
@@ IV DISCUSSION
Imagining potential benefits to speed, accuracy, and safety, we presented an overview of indoor crime scene analysis (CSA) by a nano drone:
  • Theoretical. First, we identified some theoretical gaps in eleven categories, based on a "big picture" (rapid scoping) review.
  • Practical. Second, some practical challenges, summarized in Table III, were explored by iteratively designing three prototype capabilities relating to accessing crime scenes through windows, mapping and gathering evidence, and analyzing blood smears, with performances of 75%, 85%, and 80%.
A summary video and some code were also made available [70, 71].
TABLE III: Some practical challenges identified.
Gaining entry to a crime scene through partially opened windows Low thrust force, propellers getting caught despite protective guards, and inaccurate control (downwash, contacting the window at inefficient locations near the hinges, and being knocked away).
Mapping and gathering evidence Low illumination and low image resolution, lack of contrast between objects to detect and the floor, limited amounts of training images, and drone instability
Estimating motion of bloody objects (i.e., the "dragged direction") Color picking in low illumination with standard hue values for red, detection of only one contour or many tiny contours, and amorphous stains from soaked tissues dispensing an equal amount of red at the start and end points.
In summary, while nano drones do not seem capable yet of, e.g., skillfully opening windows, gathering evidence, and analyzing complex blood patterns in the real world, we believe that our results suggest the feasibility of initial ideation and testing within a laboratory setting. Additionally, we have aimed to follow a sustainable approach–in using a low-cost, low-energy platform and documenting prototyping in an open way–toward supporting a positive societal impact.
@@ IV-A Limitations and Future Work
Our approach is limited by its exploratory nature: We expect many more important theoretical questions could exist that were not discussed here. For example, wireless transmission of potentially sensitive data from drones to remote laptops for processing could present an security threat, allowing criminals to intercept and possibly alter communications between drones and humans. As well, the scope of this paper did not allow us to explore all of the gaps we identified. Even for the explored capabilities, we expect huge discrepancies between real-world scenarios and our exploratory experiments, involving rapid, low-fidelity prototyping; e.g., using mock-up windows, small lab spaces with markers, and photos or red paint in place of real firearms or human blood.
Ideation from researchers of various backgrounds will expose new important theory. Future work should also move farther from the laboratory, and into the real world, to better make conclusions about the applicability of the designated capabilities: Stronger, more advanced drones will autonomously fly in more accurate settings, onboard sensors like lidars will be used rather than markers for mapping, and real blood stain patterns will be analyzed. In-built processing will also enhance privacy (possibly along with federated learning and selective blurring); possible disturbances (i.e., destruction of sensitive evidence) due to a drone’s static electricity, electromagnetic interference, lights, or heat should be thoroughly checked; and potential misuses by criminals identified (e.g., unlawful entry via windows for burglary, or detection and tracking of innocent citizens by stalkers). Additionally, we have started to explore a broader range of the capabilities identified as potential gaps, conducting some development on each of our ideas. These additional scenarios are visualized in Fig. 6, and in the accompanying video.
Figure 6: Additional Prototyping: (a) Revealing hidden blood stains, (b) documenting evidence without disturbance via a drifting blimp, (c) using a flashlight in the dark, and (d) detecting hidden people with thermal or radar.
In conclusion, while much work remains to be done to make this vision of nano drones helping to solve crimes a reality, we believe that the day is not too far off in which such technologies will help to preserve justice and otherwise impact society in a positive way.